You need to know the height of a room, but you have no tape measure. You fasten one end of a string to the ceiling of the room, and tie a small rock at the other end so it almost touches the floor. You start this simple pendulum swinging slightly, and measure its period, which is 3.50 seconds. How tall is the room? Answer in meters.
1
Expert's answer
2011-02-24T05:36:38-0500
The period and the length of mathematical pendulum are related by the following expression:
T = 2 π √(L/g)
Thus the length is L = T2g/4π2 = 3.52* 9.8 / (4*3.142) = 3.04 [m]